OpenStudy (anonymous):

how many times does 2/3 fit into 3 and 2/3

OpenStudy (anonymous):

Lets say x 2/3 fits into 3 \[x \times \frac{2}{3} = 3\] \[ x = \frac{3}{2} \times 3\] \[x = \frac{9}{2}\] \[x=4.5\]

OpenStudy (anonymous):

in fraction form though....

OpenStudy (anonymous):

Just keep it 9/2